If the revenue function for a noodle bowl stand is r(x)= 33/16x-42 and the cost function is c(x)=33/32x+47, how much profit (or less) will be generated each hour when 95 noodle bowls per hour are produced and sold?

Thanks so much for any help. I can't figure out how to set it up to get to the answer.
Thanks for any help you can provide.

Answer by stanbon(75887)

If the revenue function for a noodle bowl stand is r(x)= 33/16x-42 and the cost function is c(x)=33/32x+47, how much profit (or less) will be generated each hour when 95 noodle bowls per hour are produced and sold?
-------------
Profit = Revenue - Cost
P(x) = [(33/16)x-42] - [(33/32)x+47]
P(x) = (33/32)x-89
Then:
P(95)= (33/32)(95)-89
P(95)= $8.97
Cheers,
Stan H.